# Feasibility of Learning Weighted Automata on a Semiring

### v4

@@ 5. Conclusion and open questions
We have pictured a landscape for learning weighted automata on any semiring and shown that an approach à la Angluin can only work in general on a restricted class of functions. This paper opens many questions, both general and specific. First, for functions that are not weakly guessable, one can ask whether there is still a way to find appropriate states (not from rows of the Hankel matrix) by strengthening the queries. Next, for weakly guessable functions, one can ask whether there is a way to design strategies to update $Q$ and $T$ (other than closure and adding suffixes of counter-examples to $T$) that would ensure to find a suitable hypothesis automaton. In particular, is there a way to choose “cleverly” the coefficients of the linear combinations under consideration? For example, the algorithm of [36] can fail for functions that are not guessable (i.e. if for some finite set $Q$ at every Test step the counter example provided by the teacher allows for the possibility that the closure strategy produces a solution that is not in $\Lambda_{Q}$). For guessable functions, a naive way construct an algorithm that terminates with success is to use the update strategy suggested by the proof of Proposition 3.18. But is there a better way? And if so, it would be interesting to see if the class of functions learnable by a polynomial time algorithm can be characterised. Regarding the general hierarchy of functions, there is a class for which we neither have an example of non-emptiness or an argument for collapse: the functions that are guessable but not strongly guessable (and its dual). Finally, it would be interesting to study the notion of uniform strong guessability introduced in Remark 4.2 and in particular, try to characterise the class of functions lying in the intersection of the six classes we have introduced. It would also be of interest to understand how termination of the algorithm from [36] relates to strong guessability.
@@ Appendix A Dual approach using columns
If $Q=\Sigma^{*}$, we will simply write $\Gamma_{T}$ instead of $\Gamma_{Q,T}$. Moreover, if $R\subseteq\Sigma^{*}$, then to reduce notation we will also frequently write simply $\Gamma_{Q,R}$ and $\Gamma_{R}$ instead of $\Gamma_{Q,T}$ and $\Gamma_{T}$ where $T=\{[t]:t\in R\}$.
It is clear from the definitions above that $\Gamma_{Q,T}$ is non-empty if and only if $[\varepsilon]_{Q}$ is right-generated by $T$ restricted to $Q$ and $T$ is right-closed on $Q$. Also, if $Q^{\prime}\subseteq Q$, then $\Gamma_{Q^{\prime},T}\supseteq\Gamma_{Q,T}$. However, if $T^{\prime}$ is a finite set contained in or contained by $T$, there is no containment between $\Gamma_{Q,T}$ (which lies in $S^{T\cup(T\times\Sigma\times T)}$) and $\Gamma_{Q,T^{\prime}}$ (which lies in $S^{T^{\prime}\cup(T^{\prime}\times\Sigma\times T^{\prime})}$).
Definition A.2.
Let $f:\Sigma^{*}\rightarrow S$ be a function and suppose that $T$ is a finite subset of $S^{\Sigma^{*}}$ and $Q$ is a subset of $\Sigma^{*}$ such that $\Gamma_{Q,T}\neq\emptyset$. Then for each $\gamma\in\Gamma_{Q,T}$, we define a finite weighted automaton $\mathcal{H}^{co}_{Q,T,\underline{\gamma}}$ with:
@@
Here we record the definitions used in the statement
Definition A.4.
A function $f:\Sigma^{*}\rightarrow S$ is said to be weakly co-guessable if there exists a finite set $T$ of columns of the Hankel matrix of $f$ such that $\Gamma_{T}$ is non-empty.
Definition A.5.
@@ Appendix B Examples
Let $S$ be one of $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$, or $\mathbb{R}_{\max}$, and consider the function $f_{1}$ computed by the automaton depicted in Figure 1(b) and $F$ its Hankel matrix over $S$. Thus $f_{1}(w)$ is equal to the length of the longest block of consecutives $a$’s in $w$. Let $w=ua^{n}$ and $w^{\prime}=a^{n^{\prime}}v$ where $u,v\in\Sigma^{*}$ are such that $u$ does not end with an $a$ and $v$ does not start with an $a$. The corresponding entry of the Hankel matrix is given by $F_{w,w^{\prime}}=f(ww^{\prime})={\max}\{f_{1}(u),n+n^{\prime},f_{1}(v)\}$. We claim that $f_{1}$ is neither weakly guessable nor weakly co-guessable. Suppose for contradiction that $f_{1}$ is weakly guessable. By Theorem 3.6 there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the rows of the Hankel matrix indexed by $W$ are row-generating. Let us write $w_{i}=u_{i}a^{n_{i}}$ for $i=1,\ldots,k$, where $u_{i}$ does not end with an $a$ and $n_{i}\geq 0$, and choose a positive integer $N$ such that $N\geq{\max}_{i\in\{1,\ldots,k\}}\{f_{1}(u_{i}),n_{i}\}$. Consider the row of the Hankel matrix indexed by $w=a^{N+1}b$. Since the rows indexed by $W$ are assumed to be row-generating, we have: $\langle w\rangle=\bigoplus_{i=1}^{k}\lambda_{i}\otimes\langle w_{i}\rangle$ for some $\lambda_{i}\in S$. Now set $u=ba^{N}$ and $v=ba^{N+1}$. We have:
@@
Finally, $f_{2}$ is not guessable. Suppose for contradiction that $f_{2}$ is guessable. Then there exist finite sets $Q$, $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$. and hence for all $\underline{\lambda}$ in $\Lambda_{Q,T}$, we have that $\mathcal{H}_{Q,T,\lambda}$ computes $f_{2}$. By the observations at the start of this subsection, we must have that $Q$ contains $\varepsilon$ and at least one word, say $q$, beginning with $a$ and at least one word, say $q^{\prime}$, beginning with $b$. Without loss of generality, we may assume that for a fixed constant $N$ (to be chosen below) we have that the length of the longest word in $T$ is greater than $N$ (since for all $T^{\prime}\supseteq T$ we also have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}=\Lambda_{Q}\neq\emptyset$). Now let $n$ be equal to the maximal sum of the length a word in $Q$ and the length of a word in $T$ (which by the above observation can be made arbitrarily large, by modifying $T$ if necessary). For this choice of $n$ and for all $t\in T$ we have
@@
Now let $\mathcal{H_{Q,T,\underline{\lambda}}}$ be the hypothesis automaton constructed from an element $\underline{\lambda}\in\Lambda_{Q,T}=\Lambda_{Q}$ using the linear combinations above (containing the negative coefficients $-n$) and let $\mathcal{H}_{q}$ denote the automaton obtained from $\mathcal{H}_{Q,T,\underline{\lambda}}$ by making $\langle q\rangle$ the unique initial state. Since $\underline{\lambda}\in\Lambda_{Q}$, it follows from the first part of the proof of Theorem 2.12 that for all words $w$, and all $q\in Q$, $\langle q\rangle_{w}$ is equal to the value computed by $\mathcal{H}_{q}$ on input $w$. Thus in particular we have that $f_{2}(qa^{n}bb^{2n})$ is equal to the value computed by $\mathcal{H}_{q}$ on input $a^{n}bb^{2n}$. Consider the run composed of: the transition from $q$ to $q$ labelled by $a$ for each instance of $a$ in the factor $a^{n}$, the transition from $q$ to $q^{\prime}$ labelled by $b$ for the first factor of $b$, and the transition from $q^{\prime}$ to $q^{\prime}$ labelled by $b$ for each instance of $b$ in the remaining factor $b^{2n}$. Since the final weight at state $q^{\prime}$ is (by definition) $\langle q^{\prime}\rangle_{\varepsilon}=f_{2}(q^{\prime})=|q^{\prime}|_{b}$, it then follows from the equations above (recalling that $\otimes$ denotes usual addition in the semirings under consideration) that this run has weight $n+(-n)+n+|q^{\prime}|_{b}=2n+|q^{\prime}|_{b}$ giving $f_{2}(qa^{n}bb^{2n})>2n$ (recalling that the addition of this semiring is maximisation). On the other hand, since $q$ starts with the letter $a$, we have $f_{2}(qa^{n}bb^{2n})=|q|_{a}+n$ (from the definition of this function). But now taking $N=|q|_{a}$ we have $f_{2}(qa^{n}bb^{2n})>2n>n+N=|q|_{a}+n=f_{2}(qa^{n}bb^{2n})$, giving a contradiction.
@@
Since $\mathbb{Z}_{\max}$ and $\mathbb{R}_{\max}$ are commutative, and $f_{2}$ is weakly guessable but neither guessable nor weakly co-guessable, it follows from Lemma 4.3 that $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable).
@@
We will give examples of functions that are strongly guessable but not weakly co-guessable for $\mathbb{N}_{\max}$, $\mathbb{N}$, $\mathcal{P}_{\rm fin}(\Sigma^{*})$ and $\mathbb{R}_{\geq 0}$. In Lemma 4.4, it is shown that all the functions from the first three semirings are strongly row-bound and hence, any weakly guessable function is strongly guessable. It is also the case for $\mathbb{R}_{\geq 0}$ by Lemma 4.5. To prove that the functions under consideration are strongly guessable, it is then enough to prove that they are weakly guessable, or equivalently that they are computed by a literal automaton.
For $S=\mathbb{N}_{\max}$, consider the function $f_{3}$ computed by the automaton depicted in Figure 1(a) and $F$ its Hankel matrix. Thus $f_{3}(w)=|w|_{a}$ if $w$ starts with an $a$, $f_{3}(w)=|w|_{b}$ if $w$ starts with an $b$, and $f_{3}(\varepsilon)=-\infty$, and (by the same argument as for $f_{2}$) we find that $f_{3}$ is computed by a literal automaton and weakly guessable; in this case however Lemma 4.4 applies to give that $f_{3}$ is also strongly guessable (and note that the proof that $f_{2}$ is not guessable made use of negative coefficents that we do not have access to here!)) Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f_{3}(aw)=|aw|_{a}$, and $f_{3}(bw)=|bw|_{b}$. Suppose for contradiction that $f_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=|w_{i}|_{a}+1$ and $F_{b,w_{i}}=|w_{i}|_{b}+1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=N+1$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
Since $\mathbb{N}_{\max}$ and $\mathbb{N}$ are commutative, and $f_{3}$ and $f^{\prime}_{3}$ are strongly guessable but not weakly co-guessable, then (by Lemma 4.3) $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable .

### v5

@@ 5. Conclusion and open questions
We have pictured a landscape for learning weighted automata on any semiring and shown that an approach à la Angluin can only work in general on a restricted class of functions. This paper opens many questions, both general and specific. First, for functions that are not weakly guessable, one can ask whether there is still a way to find appropriate states (not from rows of the Hankel matrix) by strengthening the queries. Next, for weakly guessable functions, one can ask whether there is a way to design strategies to update $Q$ and $T$ (other than closure and adding suffixes of counter-examples to $T$) that would ensure to find a suitable hypothesis automaton. In particular, is there a way to choose “cleverly” the coefficients of the linear combinations under consideration? For example, the algorithm of [36] can fail for functions that are not guessable (i.e. if for some finite set $Q$ at every Test step the counter example provided by the teacher allows for the possibility that the closure strategy produces a solution that is not in $\Lambda_{Q}$). For guessable functions, a naive way to construct an algorithm that terminates with success is to use the update strategy suggested by the proof of Proposition 22. But is there a better way? And if so, it would be interesting to see if the class of functions learnable by a polynomial time algorithm can be characterised. Regarding the general hierarchy of functions, there is a class for which we neither have an example of non-emptiness or an argument for collapse: the functions that are guessable but not strongly guessable (and its dual). Finally, it would be interesting to study the notion of uniform strong guessability introduced in Remark 31 and in particular, try to characterise the class of functions lying in the intersection of the six classes we have introduced. It would also be of interest to understand how termination of the algorithm from [36] relates to strong guessability.
@@ Appendix A Dual approach using columns
If $Q=\Sigma^{*}$, we will simply write $\Gamma_{T}$ instead of $\Gamma_{Q,T}$. Moreover, if $R\subseteq\Sigma^{*}$, then to reduce notation we will also frequently write simply $\Gamma_{Q,R}$ and $\Gamma_{R}$ instead of $\Gamma_{Q,T}$ and $\Gamma_{T}$ where $T=\{[t]:t\in R\}$. It is clear from the definitions above that $\Gamma_{Q,T}$ is non-empty if and only if $[\varepsilon]_{Q}$ is right-generated by $T$ restricted to $Q$ and $T$ is right-closed on $Q$. Also, if $Q^{\prime}\subseteq Q$, then $\Gamma_{Q^{\prime},T}\supseteq\Gamma_{Q,T}$. However, if $T^{\prime}$ is a finite set contained in or contained by $T$, there is no containment between $\Gamma_{Q,T}$ (which lies in $S^{T\cup(T\times\Sigma\times T)}$) and $\Gamma_{Q,T^{\prime}}$ (which lies in $S^{T^{\prime}\cup(T^{\prime}\times\Sigma\times T^{\prime})}$). {defi} Let $f:\Sigma^{*}\rightarrow S$ be a function and suppose that $T$ is a finite subset of $S^{\Sigma^{*}}$ and $Q$ is a subset of $\Sigma^{*}$ such that $\Gamma_{Q,T}\neq\emptyset$. Then for each $\gamma\in\Gamma_{Q,T}$, we define a finite weighted automaton $\mathcal{H}^{co}_{Q,T,\underline{\gamma}}$ with:
@@
Here we record the definitions used in the statement {defi} A function $f:\Sigma^{*}\rightarrow S$ is said to be weakly co-guessable if there exists a finite set $T$ of columns of the Hankel matrix of $f$ such that $\Gamma_{T}$ is non-empty.
{defi}
@@ Appendix B Examples
Let $S$ be one of $\mathbb{N}_{\max}$, $\mathbb{Z}_{\max}$, or $\mathbb{R}_{\max}$, and consider the function $f_{1}$ computed by the automaton depicted in Figure 1(b) and $F$ its Hankel matrix over $S$. Thus $f_{1}(w)$ is equal to the length of the longest block of consecutives $a$’s in $w$. Let $w=ua^{n}$ and $w^{\prime}=a^{n^{\prime}}v$ where $u,v\in\Sigma^{*}$ are such that $u$ does not end with an $a$ and $v$ does not start with an $a$. The corresponding entry of the Hankel matrix is given by $F_{w,w^{\prime}}=f(ww^{\prime})={\max}\{f_{1}(u),n+n^{\prime},f_{1}(v)\}$. We claim that $f_{1}$ is neither weakly guessable nor weakly co-guessable. Suppose for contradiction that $f_{1}$ is weakly guessable. By Theorem 12 there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the rows of the Hankel matrix indexed by $W$ are row-generating. Let us write $w_{i}=u_{i}a^{n_{i}}$ for $i=1,\ldots,k$, where $u_{i}$ does not end with an $a$ and $n_{i}\geq 0$, and choose a positive integer $N$ such that $N\geq{\max}_{i\in\{1,\ldots,k\}}\{f_{1}(u_{i}),n_{i}\}$. Consider the row of the Hankel matrix indexed by $w=a^{N+1}b$. Since the rows indexed by $W$ are assumed to be row-generating, we have: $\langle w\rangle=\bigoplus_{i=1}^{k}\lambda_{i}\otimes\langle w_{i}\rangle$ for some $\lambda_{i}\in S$. Now set $u=ba^{N}$ and $v=ba^{N+1}$. We have:
@@
Finally, $f_{2}$ is not guessable. Suppose for contradiction that $f_{2}$ is guessable. Then there exist finite sets $Q$, $T$ such that $\Lambda_{Q,T}=\Lambda_{Q}\neq\emptyset$, and hence for all $\underline{\lambda}$ in $\Lambda_{Q,T}$, we have that $\mathcal{H}_{Q,T,\lambda}$ computes $f_{2}$. By the observations at the start of this subsection, we must have that $Q$ contains $\varepsilon$ and at least one word, say $q$, beginning with $a$ and at least one word, say $q^{\prime}$, beginning with $b$. Without loss of generality, we may assume that for a fixed constant $N$ (to be chosen below) we have that the length of the longest word in $T$ is greater than $N$ (since for all $T^{\prime}\supseteq T$ we also have $\Lambda_{Q,T}=\Lambda_{Q,T^{\prime}}=\Lambda_{Q}\neq\emptyset$). Now let $n$ be equal to the maximal sum of the length a word in $Q$ and the length of a word in $T$ (which by the above observation can be made arbitrarily large, by modifying $T$ if necessary). For this choice of $n$ and for all $t\in T$ we have
@@
Now let $\mathcal{H_{Q,T,\underline{\lambda}}}$ be the hypothesis automaton constructed from an element $\underline{\lambda}\in\Lambda_{Q,T}=\Lambda_{Q}$ using the linear combinations above (containing the negative coefficients $-n$) and let $\mathcal{H}_{q}$ denote the automaton obtained from $\mathcal{H}_{Q,T,\underline{\lambda}}$ by making $\langle q\rangle$ the unique initial state. Since $\underline{\lambda}\in\Lambda_{Q}$, it follows from the first part of the proof of Theorem 5 that for all words $w$, and all $q\in Q$, $\langle q\rangle_{w}$ is equal to the value computed by $\mathcal{H}_{q}$ on input $w$. Thus in particular we have that $f_{2}(qa^{n}bb^{2n})$ is equal to the value computed by $\mathcal{H}_{q}$ on input $a^{n}bb^{2n}$. Consider the run composed of: the transition from $q$ to $q$ labelled by $a$ for each instance of $a$ in the factor $a^{n}$, the transition from $q$ to $q^{\prime}$ labelled by $b$ for the first factor of $b$, and the transition from $q^{\prime}$ to $q^{\prime}$ labelled by $b$ for each instance of $b$ in the remaining factor $b^{2n}$. Since the final weight at state $q^{\prime}$ is (by definition) $\langle q^{\prime}\rangle_{\varepsilon}=f_{2}(q^{\prime})=|q^{\prime}|_{b}$, it then follows from the equations above (recalling that $\otimes$ denotes usual addition in the semirings under consideration) that this run has weight $n+(-n)+n+|q^{\prime}|_{b}=2n+|q^{\prime}|_{b}$ giving $f_{2}(qa^{n}bb^{2n})>2n$ (recalling that the addition of this semiring is maximisation). On the other hand, since $q$ starts with the letter $a$, we have $f_{2}(qa^{n}bb^{2n})=|q|_{a}+n$ (from the definition of this function). But now taking $N=|q|_{a}$ we have $f_{2}(qa^{n}bb^{2n})>2n>n+N=|q|_{a}+n=f_{2}(qa^{n}bb^{2n})$, giving a contradiction.
@@
Since $\mathbb{Z}_{\max}$ and $\mathbb{R}_{\max}$ are commutative, and $f_{2}$ is weakly guessable but neither guessable nor weakly co-guessable, it follows from Lemma 32 that $\bar{f_{2}}$ is weakly co-guessable but neither co-guessable nor weakly guessable (since a function over a commutative semiring is (weakly) guessable if and only if its mirror is (weakly) co-guessable).
@@
We will give examples of functions that are strongly guessable but not weakly co-guessable for $\mathbb{N}_{\max}$, $\mathbb{N}$, $\mathcal{P}_{\rm fin}(\Sigma^{*})$ and $\mathbb{R}_{\geq 0}$. In Lemma 33, it is shown that all the functions from the first three semirings are strongly row-bound and hence, any weakly guessable function is strongly guessable. It is also the case for $\mathbb{R}_{\geq 0}$ by Lemma 34. To prove that the functions under consideration are strongly guessable, it is then enough to prove that they are weakly guessable, or equivalently that they are computed by a literal automaton.
For $S=\mathbb{N}_{\max}$, consider the function $f_{3}$ computed by the automaton depicted in Figure 1(a) and $F$ its Hankel matrix. Thus $f_{3}(w)=|w|_{a}$ if $w$ starts with an $a$, $f_{3}(w)=|w|_{b}$ if $w$ starts with an $b$, and $f_{3}(\varepsilon)=-\infty$, and (by the same argument as for $f_{2}$) we find that $f_{3}$ is computed by a literal automaton and weakly guessable; in this case however Lemma 33 applies to give that $f_{3}$ is also strongly guessable (and note that the proof that $f_{2}$ is not guessable made use of negative coefficents that we do not have access to here!)) Let us prove now that it is not weakly co-guessable. For all words $w$, we have $f_{3}(aw)=|aw|_{a}$, and $f_{3}(bw)=|bw|_{b}$. Suppose for contradiction that $f_{3}$ is weakly co-guessable. Then there exists a finite set of words $W=\{w_{1},\ldots,w_{k}\}$ such that the columns of the Hankel matrix labelled by $W$ are column-generating and let $N>\max_{i}|w_{i}|_{a}+1$. Since for all $i$, we have $F_{a,w_{i}}=|w_{i}|_{a}+1$ and $F_{b,w_{i}}=|w_{i}|_{b}+1$, it is easy to see that the column indexed by $w=a^{N}$ (for which $F_{a,w}=N+1$ and $F_{b,w}=1$) cannot be expressed as a right-linear combination of the columns indexed by $W$.
@@
Since $\mathbb{N}_{\max}$ and $\mathbb{N}$ are commutative, and $f_{3}$ and $f^{\prime}_{3}$ are strongly guessable but not weakly co-guessable, then (by Lemma 32) $\bar{f_{3}}$ and $\bar{f^{\prime}_{3}}$ are strongly co-guessable but not weakly guessable .